Some the following DEs using Laplace Transform. 1) <3> y - 2 y = a* (0) = | (<4> Y + 4y + 4y = 0; Y(0) = 2 ; Y6u =-| - <s>
Answer #1

(3) Given differential equation y- 2y = ebt ė yro) = 1 First we take the Laplace thansform of all the terms in the diffenent23+7 47 Giren differential equation .. +491+4y=0, yco)=2, gro)- Laking Laplace transform of all the terms and plugging the i57 y + 4y = 800198 , 910) = 0, y (O=4. Taking Laplace transform of all the terms and plugging into the initial conditions,